Hello. Could someone please show me what the steps are to solve these problems ?
Solve for X
X^2 – 12 = -4x
X^2 – 81 = 0

Thank you so much.

Answer by mananth(16949)

Solve for X
X^2 – 12 = -4x
add +4x
x^2+4x-12=0
x^2+6x-2x-12=0
x(x+6)-2(x+6)=0
(x+6)(x-2)=0
x=-6 OR x=2
...

X^2 – 81 = 0
(x)^2-(9)^2
difference of two squares formula
a^2-b^2=(a+b)(a-b)
(x+9)(x-9)=0
x=+/- 9
...
m.ananth@hotmail.ca